ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 76 Ramollo, Motlatsi Pius National University of Lesotho, Roma180, LESOTHO. E-Mail: MP. Ramollo@nul. ls On the motion of a bore across a barrier^{∗} A study is carried out of the bore as it encounters a narrow ridge-type barrier in quiescent water. Both the critical and non-critical cases are considered. Expressions for surface disturbance caused are derived and the evolution equations are solved numerically.

* Published in Advances in Fluid Mechanics, 21,125-135, 1998